HETAV PATEL

Grade 11,

A shopkeeper sells three types of flower seeds A1, A2 and A3. They are sold as a mixture where the proportions are 4:4:2 respectively. The germination rates of three types of seeds are 45%, 60% and 35%. Calculate the probability. a) of a randomly chosen seed to germinate. b) that it is of the type A2, given that a randomly chosen seed does not

A shopkeeper sells three types of flower seeds A1, A2 and A3. They are sold as a
mixture where the proportions are 4:4:2 respectively. The germination rates of
three types of seeds are 45%, 60% and 35%. Calculate the probability.
a) of a randomly chosen seed to germinate.
b) that it is of the type A2, given that a randomly chosen seed does not

Grade:upto college level

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
let E1 : randomly selected seed is A1 type P(E1) = 4/10
E2 : randomly selected seed is A2 type P(E2) = 4/10
E3 : randomly selected seed is A3 type P (E3) = 2/10
(i) let A : selected seed germinates
P (A/E1) = 45/100
P (A/E2) = 60/100
P (A/E3) = 35/100
P(A) = P(E1) P(A/E1) + P(E2) P(A/E2) + P (E3). P(A/E3)
= 0.49
(ii) let A : selected seed does not germinate
P(A/E1) = 55/100
,P(A/E2) = 40/100
P(A/E3) = 65/100
P(E2/A) = ((P(E2) P(A|E2))/ P(E1)P(A|E1)+P(E2)P(A|E2)+P(E3)P(A|E3)
= 16/51
Thanks and Regards
Shaik Aasif
askIITians faculty